Mauro Products distributes a single product, a woven basket whose selling price is $22 per unit and whose variable expense is $17 per unit. The company’s monthly fixed expense is $5,500.

Required:

1. Calculate the company’s break-even point in unit sales.

2. Calculate the company’s break-even point in dollar sales. (Do not round intermediate calculations.)

3. If the company's fixed expenses increase by $600, what would become the new break-even point in unit sales? In dollar sales?
Answer #1

Contribution margin=Sales-Variable cost

=(22-17)=$5 per unit

a.Breakeven point=Fixed expense/Contribution margin

=(5500/5)=1100 units

b.Breakeven value=(1100*22)=$24200

c.New fixed expense=(5500+600)=$6100

Hence new breakeven=(6100/5)=1220 units

=(1220*22)=$26840